OpenStudy (anonymous):

Find two triangles for which A=51 degrees, a = 6.5 ft, and b = 7.9 ft

OpenStudy (dumbcow):

i would use law of sines here: \[\frac{\sin B}{7.9} = \frac{\sin 51}{6.5}\] B = 70.83 or 109.17

OpenStudy (anonymous):

yay! that's what i got :)

OpenStudy (anonymous):

I also found C = 58 degrees but in the book there also another answer to C' which = 20 degrees, I don't get how they got 20 degrees ^^!

OpenStudy (anonymous):

I understand B' = 180 - 71 = 109 degrees tho :)

OpenStudy (dumbcow):

sum of angles is 180 C = 180 - 51 - 70.83 = 58 C = 180 - 51 - 109.17 = 20

OpenStudy (anonymous):

ohhh i see